The​ half-life of a drug in the bloodstream is 18 hours. What fraction of the original drug dose remains in 48 ​hours? in 60 ​hours?
Answer #1

Fraction of original drug that remains in 48 hours = (1/2)48/18 = (1/2)8/3 = 0.1575

Fraction of original drug that remains in 60 hours = (1/2)60/18 = (1/2)10/3 = 0.0992